### v1

@@ 3.4 Overall system performance
Figure 12: Transfer function of a GRAND detection unit. The transfer function is simulated using GRANDlib [48]. Left: In GP300, with a 20-dB gain at the FEB. (Differences between this figure and Fig. 8 in Ref. [48] are due to a revision of the S-parameters in the RF chain currently used by GRANDlib.) Right: In G@A, with a 0-dB gain at the FEB. For G@A, the transfer function of the X arm is exactly superimposed on the Y arm. The differences between the transfer functions of GP300 and G@A echo the differences in the setup and configuration between the arrays. For both arrays, the system response has a strong frequency dependence. See Section 3.4 for details.
@@
The complete system performance is simulated via the implementation of the radio-frequency chain (RF chain) in GRANDlib [48], a software tool developed and tailored specifically to the needs of the GRAND Collaboration. The RF chain receives as input an open-circuit voltage, $V_{\rm oc}$, which emulates the output voltage from the antenna arms. The final output voltage of GRANDlib, $V_{\rm out}$, is the resulting signal after passing through the matching network, LNA, cables, connectors, and FEB. Reference [48] contains more details of GRANDlib, the RF chain, and its implementation.
The implementation of the RF chain in GRANDlib is modular and can be adapted to the different setups of GP300 and G@A. The RF chain is described via scattering parameters (S-parameters), which quantify the reflection and transmission of the signals as they travel through the different components of a GRAND DU. The S-parameters of GP300 and G@A were measured separately in the laboratory and fed to GRANDlib.
Figure 12 shows the resulting simulated overall system response as a function of frequency for GP300 and G@A. The response is quantified via the transfer function, defined as the ratio $V_{\rm out}/V_{\rm oc}$, measured with respect to the base of the DU (see also Fig. 8 in Ref. [48]). Within GP300 and G@A, the transfer functions of the X and Y arms are similar to each other and different from the transfer function of the Z arm, reflecting their VELs (Fig. 6). However, the transfer functions are different in GP300 and G@A, primarily because of their different LNAs. Although refinement of the RF chain is still ongoing, the results in Fig. 12 already allow us to validate the setups by comparing simulations to measurements of, for example, the Galactic radio background [29].
3.5 Communications and power supply
Communication between the local and central DAQ software is done through TCP/IP sockets for both G@A and GP300. Upon request from the central DAQ, the complete event is sent over the wireless network. This network is created through M5 (G@A) or AirMax (GP300) ™Ubiquiti Bullets, Rockets, and a parabolic communications antenna. For G@A these were provided by the Pierre Auger Collaboration. Bi-directional transmission tests were performed point-to-point, i.e., DU to DAQ, in both arrays, using a fixed file size. The uplink speed, from DU to DAQ, was on average 2.9 MB s-1 for GP300 and 2.1 MB s-1 for G@A. The downlink speed, from DAQ to DU, was on average 4.2 MB s-1 for GP300 and 1.1 MB s-1 for G@A. The local and central DAQ software pipeline is described in the next section.
@@ 3.5 Communications and power supply
In both arrays, the power supply of each DU consists of a solar panel, a charge controller, and a battery. The power consumption of a single DU is roughly 15 W. The power supply components were specified to ensure the DUs can operate for one week on a full battery, even without input from the solar panel. In GP300, the solar panels have a nominal power of 180 W, with a maximum power voltage of 19.5 V. Low-noise charge controllers (Genasun GV-10-Pb-12V) were chosen to efficiently charge the battery from the solar panel as well as provide power to the FEBs. Cold-resistant gel batteries were selected, operating at a nominal output voltage of 12 V with a capacity of 200 Ah.
@@
In G@A, the AERA electronics compartment and power harvesting were mostly re-used, as the power requirements of the prototype GRAND FEB match the AERA electronics quite closely. The solar panel was the same originally deployed in the AERA-II stations (Kyocera 135 W), with a nominal power of 135 W and maximum output voltage of 16 V. New charge controllers, of the same model as deployed at GP300, were installed to replace the original AERA ones. The batteries were also exchanged for lead-acid batteries with a nominal output voltage of 12 V and 180 Ah capacity. Overall, the final power supply for G@A has a smaller capacity than for GP300. As a result, G@A DUs may need to hibernate during Argentinian winters, depending on their battery levels.
@@
In GP300 and G@A, the charge controllers are housed in Faraday cages. To further isolate them from electromagnetic noise, a filter (Spectrum Control 1250-054) was added to the output of the charge controller. The Bullet is also housed in the FEB box at G@A, and, due to the heat-dissipation modifications, in a separate Faraday cage at GP300.
4 Self-trigger and data acquisition
4.1 Self-trigger algorithm
@@
The main science case of GRAND—the study of ultra-high-energy particles [1]—involves the detection of extensive air showers. The radio emission from air showers induces a temporary distortion of the local electric field, with a duration of several nanoseconds. Through the data acquisition chain described above, this distortion is translated into a short transient in the voltage input to the ADC. However, anthropogenic sources also emit radio pulses with similar characteristics. Thus, air-shower experiments that rely solely on their radio-detection, like the GRAND prototype arrays, must have a self-triggering procedure that selects only detected events that resemble those expected from air showers.
@@
Figure 13: Local Trigger flow for an incoming signal at a GRAND detection unit. The trigger algorithm starts when the amplitude of the signal exceeds the preset “signal threshold", i.e., when a signal crossing happens, as long as there was no such crossing in the preceding quiet time window. In the time period following the signal crossing, called the “trigger period window”, the algorithm counts the number of crossings of a lower “noise threshold". If the number of noise-threshold crossings is between predefined allowed minimum and maximum values, and the time interval between any two adjacent noise-threshold crossings does not exceed a predefined maximum value, the trigger process is completed and the signal is accepted. Because the trigger logic runs under the 500-MHz clock in the FPGA, the timing has a resolution of 2 ns. The time stamp, corresponding to the time of the end of the trigger period window, is recorded and it is sent to the Central DAQ.
The detected radio pulse emitted by an air shower generates a short-duration, high-amplitude (relative to the average baseline) oscillation in the FEB output. Man-made noise usually creates a wider pulse or a train of pulses. In addition, constructive interference of single-frequency anthropogenic sources can also create air shower-like pulses. The Local Trigger happens at the DU level and is implemented within the firmware running on the FPGA chip. Its aim is to select air-shower candidates while rejecting background pulses by efficiently identifying peaks in the difference of two consecutive ADC samples. With the prototype GRAND FEB, it is possible to first remove single-frequency noise sources with a digital filter, as shown in Section 3.3. The resulting filtered signal is fed to the trigger algorithm.
Figure 13 shows a representative flowchart of the trigger algorithm, which begins with a configurable “signal threshold”. If the signal exceeds this threshold and there has not been any threshold crossing in a time window prior to the current pulse, the algorithm starts. The signal threshold is set to a value that is suitably higher than the average noise level, while maintaining a reasonable trigger rate. The second of these conditions aims to avoid pulse trains. Next, in a time window after the pulse, the system counts how many times a second “noise threshold” is crossed.
@@
The width of the time windows, the signal and noise threshold levels, the allowed maximum number of crossings of the noise threshold, and the maximum time between successive crossings are adjustable parameters. They are set offline at the Central DAQ prior to each data acquisition run. The values are chosen to best reject long and/or irregular pulse trains, and to prevent digital glitches from corrupting the voltage traces. This trigger flexibility allows us to adapt to the different noise environments of each array, with G@A requiring stricter settings to avoid the more frequent narrow and broadband radio frequency interference (RFI). When all triggering conditions are met, the Local Trigger is complete and the timestamp is sent to the Central DAQ.
@@
Next to recording air showers, the arrays also keep track of the background conditions via unbiased triggering. This corresponds to periodically recorded data, independent of the measured voltage levels. Such unbiased data are suitable for monitoring the environment and for calibrating the setups using the periodic variation of the Galactic noise, which is currently ongoing [36, 30]. Unbiased data can also be used for additional science cases like measuring fast radio bursts [48]. Unbiased triggering, by default, takes place once every 10 s (a pre-defined 20-Hz mode is also available). In the G@A firmware version, it is possible to set the unbiased triggering frequency between 1 Hz and 100 kHz. GP300 and G@A perform both transient and unbiased triggering.
@@
Achieving a high self-trigger performance is key to the success of a standalone radio-detection experiment such as GRAND. The achieved trigger rate as a function of the incoming rate of transient pulses is of particular importance. For the GP300 firmware, the trigger rate was measured using a signal source generating four periods of a 150-MHz sine wave fed into one channel of the GRAND FEB. A 100% trigger efficiency was measured up to a repetition rate of 1434 Hz, a value larger than the nominal 1 kHz set for trigger at the DU level. These results also apply to G@A, which registers trigger rates of about 1 kHz on-site without losing functionality.
@@ 4.2 Data acquisition software
Figure 14: Data acquisition (DAQ) flow for triggered events in the GRANDProto300 and GRAND@Auger. The locally triggered event at the FPGA firmware is stored by the Local DAQ, which sends the corresponding timestamps to the Central DAQ. The Central Trigger conditions are checked at the Central DAQ and, if met, the complete event information is requested of each Local DAQ. The full event is sent and erased from local memory and the Central DAQ writes the final data to disk. The cycle repeats indefinitely while the data-taking is ongoing. Unbiased triggered events bypass the Local Trigger at the FPGA but still go through the same DAQ pipeline. See Section 4 for details.
@@
Figure 14 shows a diagram of the GRAND DAQ software. The software operates on two levels—the Local DAQ at each DU and the Central DAQ at the Central Station Computer of each array—which interact to manage the complete data flow. The Collaboration developed two separate versions of the DAQ software for GP300 and G@A to assess different implementation strategies. For the next stages of GRAND, a unified and optimized version will be developed.
@@
The Local DAQ temporarily saves both types of triggered data, transient and unbiased, on local storage, used as a circular buffer. The GP300 firmware uses the Petalinux 30-MB memory, while the G@A firmware uses the two 4-GB DDR memories. Events are stored locally while the Central Trigger, running on the Central DAQ, determines if they are to be retrieved and saved permanently.
@@
The Central Trigger decision is based upon the timing and transient trigger condition information that the Local DAQ sends to the Central DAQ. The trigger decision is determined by configurable parameters such as the maximum time difference between measured signals at different DUs and the number of DUs that detect a signal in coincidence within this time window. As soon as there is a positive decision by the Central Trigger, the Central DAQ requests the full event information from the Local DAQ. The full event information is sent over the wireless network to the Central DAQ and its event-building process. This process combines all information into a single event and writes the data to the local disk. All coincident transient triggered data that satisfy the Central Trigger are recorded in coincident data (CD) binary files.
@@
Furthermore, a random selection of local transient triggered events can be promoted to a Central Trigger and be saved as well. These events have no coincidences with other DUs and are used to monitor the trigger algorithm and the noise environment. They are stored separately in unit data (UD) binary files. Unbiased triggers are always promoted to Central triggers and are saved in monitoring data (MD) files.
5 Data storage
@@
The final data files of each trigger type—CD, UD, and MD—are saved on the local storage of the Central DAQ computer in GP300 and G@A. For GP300, the binary files are sent in real time to a computer center in Nanjing, China, through a wireless bridge. The files are then manually moved to the CC-IN2P3 cluster in Lyon, France. For G@A, new files on disk are copied daily to the central storage in CC-IN2P3 using the 4G network. This copying does not interfere with the data-taking or the bandwidth used by the Pierre Auger Collaboration. In case of high RFI conditions or failed off-site data transfer, the DAQ computer has enough memory to store all data.
@@
The automatic copying requires a local bash script to run automatically (i.e., a cron job) on the central computer for G@A and semi-manually for GP300. This script first searches for files not already transferred to CC-IN2P3 and stores this list in a simple local SQLite database. This database is used to keep track of previously transferred files and to determine if a file needs to be sent. Afterwards, the new data files are relayed to the corresponding directories on CC-IN2P3 using rsync and then flagged into the local SQLite database.
@@
When data dispatch is finalized, a second script is triggered in CC-IN2P3 that registers all newly transferred files in the main database and launches the conversion into the GRANDRoot format, described in Ref. [48]. The GRANDRoot files are moved to the corresponding directories and registered in the main database.
6 Data quality monitoring

### v2

@@ 3.4 Overall hardware system performance
Figure 10: Transfer function of a GRAND detection unit, simulated using GRANDlib [13]. Left: In GP300, with a 20-dB gain at the FEB. (Differences between this figure and Fig. 8 in Ref. [13] are due to a revision of the S-parameters in the RF chain currently used by GRANDlib.) Right: In G@A, with a 0-dB gain at the FEB. For G@A, the transfer function for the North-South polarization is exactly superimposed on the East-West polarization.
@@
The complete hardware system performance is simulated via the implementation of the Radio-Frequency chain (RF chain) in GRANDlib [13], a software tool developed and tailored specifically to the needs of the GRAND Collaboration. The RF chain receives as input an open-circuit voltage, $V_{\rm oc}$, which emulates the output voltage from the antenna arms. The final output voltage of GRANDlib, $V_{\rm out}$, is the resulting signal after passing through the matching network, LNA, cables, connectors, and FEB. Reference [13] contains more details of GRANDlib, the RF chain, and its implementation, which is modular and can be adapted to the different setups of GP300 and G@A. The RF chain is described via Scattering parameters (S-parameters), which quantify the reflection and transmission of the signals as they travel through the different components of a GRAND DU. The S-parameters of GP300 and G@A were measured separately in the laboratory and fed to GRANDlib.
Figure 10 shows the resulting simulated overall hardware system response as a function of frequency for GP300 and G@A. The response is quantified through the transfer function, defined as the ratio $V_{\rm out}/V_{\rm oc}$, measured with respect to the base of the DU (see also Fig. 8 in Ref. [13]). Within GP300 and G@A, the transfer functions of the North-South and East-West polarizations are similar to each other, a feature also reflected in their VELs (Fig. 6, top). For both arrays, the system response is strongly frequency dependent.
3.5 Power supply and electromagnetic shielding
@@
In both arrays, the power supply of each DU consists of a solar panel, a charge controller, and a battery. The power consumption of a single DU is roughly 15 W. The power supply components were specified to ensure that the DUs can operate for one week on a full battery, even without input from the solar panel. In GP300, the solar panels have a nominal power of 180 W, with a maximum power voltage of 19.5 V. Low-noise charge controllers (Genasun GV-10-Pb-12V) were chosen to efficiently charge the battery from the solar panel and provide power to the FEBs. Cold-resistant gel batteries were selected, operating at a nominal output voltage of 12 V with a capacity of 200 Ah.
@@
In G@A, the AERA electronics compartment and power harvesting were mostly re-used, as the power requirements of the prototype GRAND FEB match the AERA electronics quite closely. The solar panel was the same originally installed in the AERA-II stations (Kyocera 135 W), with a nominal power of 135 W and maximum output voltage of 16 V. New charge controllers, of the same model as deployed at GP300, were installed to replace the original AERA ones. The batteries were also exchanged for lead-acid batteries with a nominal output voltage of 12 V and 180 Ah capacity. Overall, the final power supply for G@A has a smaller capacity than for GP300. As a result, G@A DUs may need to hibernate during Argentinian winters, depending on their battery levels.
@@
In GP300 and G@A, the charge controllers are housed in Faraday cages. To further isolate them from noise and RFIs, a filter (Spectrum Control 1250-054) was added to the output of the charge controller. The Bullet is also housed in the FEB box at G@A, and, due to the heat-dissipation modifications, in a separate Faraday cage at GP300.
3.6 Communications
Communication between the local and central DAQ softwares is done over TCP/IP sockets for both G@A and GP300. Upon request from the central DAQ, the complete event is sent over the wireless network, which is created through M5 (G@A) or AirMax (GP300) ™Ubiquiti Bullets and Rockets. The rocket is placed at the Central Station and each GRAND DU has a bullet connected to its FEB by an ethernet cable. The parabolic communication antenna connects to the bullet and focuses its transmission to the rocket position. The rocket is connected to a planar antenna aimed to receive signals from multiple locations. A proprietary time sharing protocol is used to enable fast and stable communication. For G@A, the bullets and rocket were provided by the Pierre Auger Collaboration.
Bi-directional transmission tests were performed point-to-point, i.e., single DU to DAQ, in both arrays, using a fixed file size. The uplink speed, from DU to DAQ, was on average 2.9 MB s-1 for GP300 and 2.1 MB s-1 for G@A. The downlink speed, from DAQ to DU, was on average 4.2 MB s-1 for GP300 and 1.1 MB s-1 for G@A.
For the overall network to the central DAQ computer, the Ubiquity system achieves a transfer rate of 9 MB s-1, which is shared between all DUs. Taking into account typical GRAND event sizes, a bandwidth of 2 KB s-1 per station is needed. Therefore, communication between the DUs and the central DAQ does not limit the total throughput of the system. However, data transfer across ethernet has some latency, so we allow up to 3 seconds to ensure that all data is received by the central DAQ. Central trigger decision making—see Section 4 for details—and transfer back can take up to an additional second, which results in a total time of 4 seconds for which data needs to be buffered locally. Thus, the transfer from FGPA to DDR memory in the DU FEB limits the total transfer rate to roughly 1400 Hz.
4 Data acquisition processes
@@
The main science case of GRAND—the study of ultra-high-energy particles [11]—involves the radio-detection of extensive air showers. These emit electromagnetic radiation in the radio frequency range, which induces a temporary distortion of the local electric field with a duration of several nanoseconds. The radio-frequency pulse generates a high-amplitude (relative to the average baseline), short-duration oscillation at the GRAND DU ADC output, which allows for its detection. However, Radio Frequency Interference (RFI) pulses, present in the detector array region, can have similar characteristics. The main differences to air shower-induced pulses are larger low frequency contributions and/or arriving in a train of pulses. Thus, air shower experiments that rely solely on their radio-detection, like the GRAND prototype arrays, must have a data acquisition chain with a self-triggering procedure that efficiently selects detected events which most resemble those expected from air showers.
4.1 Firmware and self-trigger algorithm
The lowest level of data handling is carried out by the firmware running on the FPGA and CPU chip—specifications are given in Section 3.3. The firmware continuously retrieves the digital data from the ADC. It is configured such that, for each input sample, the previous 32 (64 ns) are used to evaluate the true baseline and perform a dynamic baseline subtraction. Sixteen independent second-order digital Infinite Impulse Response (IIR) notch filters [51], four per signal input, can be used to digitally shape the signals. Each filter has a pole frequency and width that can be individually configured. They allow us to first remove single-frequency background and then feed the resulting time-traces to the trigger algorithm. Figure 11 shows the performance of the digital filter using three independent notch filters on data collected in the G@A prototype array. If signal shaping is selected, the shaped signals are used for self-triggering. Otherwise, raw data are used. For the final readout, it is possible to select raw or shaped data.
Figure 11: Measured performance of the digital notch filters on the prototype GRAND FEB. The filter pole frequencies were set at 95.5, 151.5, and 207.7 MHz (in pink), and were applied to North-South polarization data. We show the ratio of the filtered and the unfiltered data (in blue), in which the decrease in intensity at the pole frequencies demonstrates proper filtering performance.
@@ 4.1 Firmware and self-trigger algorithm
The firmware performs the Local Trigger, at the DU level. The aim is to select air-shower candidates while rejecting background RFI pulses by efficiently identifying peaks using the difference of two consecutive ADC samples. Figure 12 shows a representative flowchart of the trigger algorithm. It starts when the amplitude of the input signal exceeds the preset “signal threshold", i.e., when a signal threshold crossing occurs. This condition is valid as long as there was no such crossing in the preceding quiet time window. In the time period following the signal crossing, called the “trigger period window”, the algorithm counts the number of crossings of a lower “noise threshold". If the number of noise-threshold crossings is between a predefined allowed minimum and maximum values, and the time interval between any two adjacent noise-threshold crossings does not exceed a predefined maximum value, the trigger process is completed and the signal is accepted.
Figure 12: Local Trigger flow for an incoming signal at a GRAND detection unit. The signal depicted represents a typical air shower-induced electric field pulse in a limited frequency band. The configurable trigger parameters are indicated w.r.t. the radio-frequency pulse, with time increasing to right.
@@
The width of the time windows, the signal and noise threshold levels, the allowed maximum number of crossings of the noise threshold, the maximum time between successive crossings, and the time-series length are adjustable parameters within the firmware. They are set by the Central DAQ prior to each data acquisition run. The values are chosen to best reject long signals and/or pulse trains and to prevent digital glitches from corrupting the voltage traces. This trigger flexibility allows us to adapt to the different noise and background environments of each array.
Although many trigger configurations were tested during the commissioning period of both prototype arrays, there was a common convergence. The signal threshold value is usually set to five times the standard deviation of the typical time-series noise level. The noise threshold is configured to be on the order of two times the amplitude of the same noise level. In GP300, keeping these threshold levels fixed is enough to ensure a reasonable trigger rate. Due to the more frequent narrow and broadband RFIs in G@A, these were used as initial threshold levels, which are dynamically adjusted by the firmware to maintain the trigger rate up to a certain value.
[…]
When the trigger conditions are met, the different information fractions that make up the event are placed in the appropriate registers. For each time-trace, two registers are used—pre and post trigger. One register is filled with the event header and the second with the GPS information. The firmware indicates when all these processes are completed, and the Local DAQ software, running under PetaLinux, initiates the transfer from the registers into memory.
Two firmware versions were tested, one on each prototype array, which differed in the method of data transfer into memory. The firmware version deployed in GP300 starts by reading event packages by the Advanced eXtensible Interface (AXI) bus, as long as the FIFO queue for storing event data is not empty. Given the quieter electromagnetic environment of GP300, the AXI speed is enough to deal with the trigger rate. However, the higher trigger rate at the G@A site required a faster memory transfer method. Therefore, the software uses Direct Memory Access (DMA transfer), moving all parts of each triggered event into the DDR memory. Hence, each fragment is moved to the correct memory position and the final event is created. The event timestamp, corresponding to the time at the end of the trigger period window, is recorded and sent to the Central DAQ software. The complete events are locally stored in the ring buffer until requested by the Central DAQ or overwritten.
@@
Next to recording air showers, the arrays also keep track of the electromagnetic environment conditions via unbiased triggering. This corresponds to periodically recorded data, independent of the measured voltage levels. The unbiased data are suitable for monitoring the local RFI background and for calibrating the DUs by using the periodic variation of the Galactic synchrotron emission—further details are provided in Section 6. Unbiased triggering, by default, takes place once every 10 seconds (a pre-defined 20-Hz mode is also available). In the G@A firmware version, it is possible to set the unbiased triggering frequency between 1 Hz and 100 kHz. GP300 and G@A perform both transient and unbiased triggering.
@@
Trigger efficiency as a function of input rate was measured with the GP300 firmware using a signal source that generated four periods of a 150-MHz sine wave fed into one channel of the GRAND FEB. A 100% trigger efficiency was measured up to a repetition rate of 1434 Hz, a value higher than the nominal 1 kHz trigger rate at DU level. These results also apply to G@A, which has registered trigger rates of about 1 kHz on-site without losing functionality. Under typical conditions, the DU trigger rate is on the order of 100 MHz for both prototype arrays. An efficient self-trigger algorithm is a main priority to achieve the GRAND science case and will be further detailed in a distinct publication.
@@ 4.2 Data acquisition software
Figure 13: Data acquisition flow for triggered events in the GRANDProto300 and GRAND@Auger. The cycle repeats indefinitely while the data-taking is ongoing. Unbiased triggered events bypass the Local Trigger at the FPGA but still go through the same DAQ pipeline.
@@
Figure 13 shows a diagram of the GRAND DAQ software. It operates on two levels—the Local DAQ at each DU and the Central DAQ at the Central Station Computer of each array—which interact to manage the complete data flow. The Collaboration developed two separate versions of the DAQ software for GP300 and G@A to assess different implementation strategies. For the next stages of GRAND, a unified and optimized version will be developed.
@@
The Local DAQ temporarily saves both types of triggered data, transient and unbiased, in local storage, used as a circular buffer, step (1) in Fig. 13. The GP300 firmware uses the Petalinux 30 MB memory, while the G@A firmware uses a 512 MB DDR memory. Events are stored locally, while the Central Trigger, running on the Central DAQ, determines if they are to be retrieved and saved permanently.
@@
The Central Trigger decision is based on the timing and transient trigger condition information that the Local DAQ sends to the Central DAQ, step (2) in Fig. 13. The trigger decision, step (3), is determined by configurable parameters, such as the maximum time difference between the measured signals at different DUs and the number of DUs that detect a signal in coincidence within this time window. As soon as there is a positive decision by the Central Trigger, the Central DAQ requests the full event information from the Local DAQ, step (4) in Fig. 13. The full event information is sent over the wireless network to the Central DAQ and its event-building process, step (5a), and the Local DAQ memory is cleared, step (5b) in Fig. 13. This process combines the requested full event information and the local trigger timestamps into a single event and writes it to the corresponding data file on the local disk of the Central DAQ, step (6) in Fig. 13.
@@
All coincident transient triggered data that satisfy the Central Trigger are recorded in Coincident Data (CD) binary files. Furthermore, a random selection of local transient triggered events can be promoted to a Central Trigger and be saved as well. These events have no coincidences with other DUs and are used to monitor the trigger algorithm and the noise environment. They are stored separately in Unit Data (UD) binary files. Unbiased triggers are always promoted to Central triggers and are saved in Monitoring Data (MD) files.
5 Data management
5.1 Storage
@@
The final data files of each trigger type—CD, UD, and MD—are saved in the local storage of the Central DAQ computer on GP300 and G@A. For GP300, the binary files are sent in real time to a computer center in Nanjing, China, through a wireless bridge. The files are then manually moved to the CC-IN2P3 cluster in Lyon, France. For G@A, new files on disk are copied daily to the central storage in CC-IN2P3 using the 4G network. This copying does not interfere with the data-taking or the bandwidth used by the Pierre Auger Collaboration. In case of high RFI conditions or failed off-site data transfer, the DAQ computer has enough disk space to store several months of triggered data.
@@
Automatic copying requires a local bash script to run automatically (i.e., a cron job) on the central computer for G@A and semi-manually for GP300. This script first searches for files not already transferred to CC-IN2P3 and stores this list in a simple local SQLite database. This database is used to keep track of previously transferred files and to determine if a file needs to be sent. Afterwards, the new data files are relayed to the corresponding directories on CC-IN2P3 using rsync and then flagged into the local SQLite database.
@@
When data dispatch is completed, a second script is triggered in CC-IN2P3 that registers all newly transferred files in the main database and launches the conversion into the GRANDRoot format, described in Ref. [13]. The GRANDRoot files are moved to the corresponding directories and registered in the main database.
5.2 Quality monitoring